# HawkBench: Investigating Resilience of RAG Methods on Stratified Information-Seeking Tasks

### v1

@@ Abstract
In real-world information-seeking scenarios, users have dynamic and diverse needs, requiring RAG systems to demonstrate adaptable resilience. To comprehensively evaluate the resilience of current RAG methods, we introduce HawkBench, a human-labeled, multi-domain benchmark designed to rigorously assess RAG performance across categorized task types. By stratifying tasks based on information-seeking behaviors, HawkBench provides a systematic evaluation of how well RAG systems adapt to diverse user needs.
Unlike existing benchmarks, which focus primarily on specific task types (mostly factoid queries) and rely on varying knowledge bases, HawkBench offers: (1) systematic task stratification to cover a broad range of query types, including both factoid and rationale queries, (2) integration of multi-domain corpora across all task types to mitigate corpus bias, and (3) rigorous annotation for high-quality evaluation.
HawkBench includes 1,600 high-quality test samples, evenly distributed across domains and task types. Using this benchmark, we evaluate representative RAG methods, analyzing their performance in terms of answer quality and response latency. Our findings highlight the need for dynamic task strategies that integrate decision-making, query interpretation, and global knowledge understanding to improve RAG generalizability. We believe HawkBench serves as a pivotal benchmark for advancing the resilience of RAG methods and their ability to achieve general-purpose information seeking. We will release our codes and data in this repository.
HawkBench: Investigating Resilience of RAG Methods on Stratified Information-Seeking Tasks
Hongjin Qian1, Zheng Liu1thanks: Corresponding author., Chao Gao 4, Yankai Wang3 Defu Lian3, Zhicheng Dou2 1 Beijing Academy of Artificial Intelligence 2 Gaoling School of Artificial Intelligence, Renmin University of China 3 University of Science and Technology of China 4 The Hong Kong University of Science and Technology {chienqhj,zhengliu1026}@gmail.com
@@ 1 Introduction
Large Language Models (LLMs) excel in general reasoning and knowledge-based tasks but often struggle with timeliness and knowledge coverage gaps, particularly in specialized domains and user-specific data OpenAI (2023); DeepSeek-AI et al. (2024). To address these limitations, incorporating external knowledge has become a common approach, with Retrieval-Augmented Generation (RAG) emerging as an effective solution to enhance factual accuracy and adaptability Zhu et al. (2024).
@@
During the information-seeking process using RAG, users may have a wide range of information needs, from simple factoid retrieval to more complex rationale-based queries Qian et al. (2024b); Zhao et al. (2024b). This versatility requires RAG systems to possess diverse capabilities, including accurate referencing and advanced reasoning skills.
@@
Recent advancements in RAG methods have enhanced vanilla RAG systems by targeting specific advanced capabilities. For instance, some methods focus on improving multi-hop reasoning to handle tasks with implicit information intents Zhao et al. (2024a); Xu et al. (2024), while others address information aggregation tasks by constructing intermediate structures, such as graphs or memory modules, to better integrate relevant information Qian et al. (2024b); Edge et al. (2024).
@@
While these advancements enable RAG systems to effectively leverage external knowledge for specific tasks, their ability to generalize across diverse scenarios remains uncertain. A recent survey categorizes external knowledge-based tasks into distinct levels, emphasizing that no single method can effectively address all query types Zhao et al. (2024b). This suggests that current RAG methods lack the resilience required for general-purpose information-seeking tasks, highlighting the need for a systematic evaluation of RAG methods across a broad range of information-seeking tasks, examining the resilience of these methods when faced with information-seeking tasks in any form.
@@
Existing public benchmarks for RAG evaluation focus narrowly on isolated dimensions of information-seeking tasks. For instance, LegalBench-RAG evaluates information-seeking tasks in the legal domain Pipitone and Alami (2024), MutiHop-RAG tests multi-hop reasoning Tang and Yang (2024), and CRAG emphasizes comprehensive evaluation on factual QA tasks Yang et al. (2024). While these benchmarks excel in their targeted domains, they collectively fail to assess the resilience of RAG methods across stratified task types due to three critical limitations:
@@
First, fragmented evaluation protocols. Current benchmarks are siloed by design, each prioritizing distinct query types. This specialization creates inconsistent evaluation criteria, hindering fair comparisons of RAG performance across diverse task categories. Second, domain bias and knowledge leakage. Many benchmarks rely on heterogeneous knowledge bases (e.g., Wikipedia and web snippets), leading to corpus-dependent performance gaps that obscure true method capabilities. Worse, LLMs are often pretrained on these same sources (e.g., Wikipedia), inflating benchmark scores through memorization rather than genuine retrieval-augmented reasoning. Third, limited query diversity. Most benchmarks disproportionately emphasize factoid questions (e.g., "When was Einstein born?"), neglecting rationale-based queries (e.g., "Explain how relativity revolutionized physics") that require synthesis and contextual analysis. This narrow focus misaligns with real-world user needs, where information-seeking behaviors span both factual lookup and complex exploratory reasoning.
In this paper, we introduce HawkBench, a human-labeled, multi-domain benchmark designed to systematically evaluate the resilience of RAG methods across stratified information-seeking tasks. Unlike existing benchmarks, HawkBench provides a structured evaluation framework that facilitates fair and comprehensive comparisons of diverse RAG approaches. HawkBench is characterized by the following key features:
@@
Systematic Task Stratification – We systematically define four query types: (1) explicit factoid queries, (2) implicit factoid queries, (3) explicit rationale queries, and (4) implicit rationale queries. This stratification, inspired by Zhao et al. (2024b) with refined modifications, ensures comprehensive task coverage. Importantly, all query types share the same underlying knowledge distribution, allowing for direct and fair performance comparisons across different tasks.
@@ 2.3 Comparison of the Four Query Levels
In Figure 1 (right), we compare the four query levels across two aspects: Reference and Reasoning.
First, in terms of Reference, the amount of relevant knowledge required increases from Level 1 to Level 4 queries, reflected in the mutual information between the knowledge base and retrieved knowledge, $I({\mathcal{X}},{\mathcal{Z}})$. Level 1 queries require minimal knowledge, as answers are directly extractable from a few text chunks. In contrast, higher-level queries, such as Level 3 and Level 4, require synthesizing information from a broader range of texts.
Second, in terms of Reasoning, complexity increases across levels due to the growing semantic gap between retrieved knowledge and the final answer. For Level 1 queries, reasoning is minimal, but for Level 3 and Level 4 queries, more reasoning is needed to connect multiple, loosely connected pieces of information. This is reflected in the decreasing mutual information $I({\mathcal{Z}},{\mathcal{Y}})$ as redundant information is filtered out during refinement.
@@ 2.4 Construction
While most current LLMs are proficient in general world knowledge due to their training on large-scale corpora, they often lack coverage in specialized, domain-specific areas. To address this gap, HawkBench incorporates 229 domain-specific texts into its knowledge base. These texts cover a wide range of domains, including professional textbooks (manually labeled into categories such as technology, humanities, art, and science), as well as financial reports, legal contracts, novels, and academic papers. This diverse and comprehensive collection ensures that HawkBench can thoroughly evaluate the domain resilience of RAG methods by covering a broad range of user information needs.
@@
(2) Question-Answer Pair Generation: The system prompts the LLM agent using built-in QA generation prompts to produce initial question-answer pairs. During this step, the system first samples from the knowledge base, selecting a random text span of varying lengths based on the task type. For Level-1 tasks, approximately 1K tokens are used as the context. For Level-2 tasks, we use a retrieval system retrieves the top-10 passages based on the generated L1 query, selecting five passages to prompt the agent to transform explicit factoid queries into implicit intent queries. For Level-3 and Level-4 tasks, up to 120K tokens are sampled as the knowledge context to guide the agent in generating information aggregation queries, with different prompts controlling the process. The codes for annotation system and all built-in prompts are released in this repository.
Level Discard % Edit % Ave. Time Total Time
@@
Table 1: Statistical Details of HawkBench Construction.
@@
3 Experiment
3.1 Baselines and Metrics
To investigate the resilience of RAG methods on HawkBench, we select the following representative baseline methods: Vanilla RAG: This method retrieves the top passages as context. Enhanced RAG Methods: HyDE Gao et al. (2023) generates a hypothetical document to enhance query retrieval. RQRAG Chan et al. (2024) rewrites the input query into sub-queries to refine retrieval. Global RAG: These methods index the knowledge base into an intermediate form to enhance global awareness. This includes memory-based methods such as MemoRAG Qian et al. (2024b) and graph-based methods like GraphRAG Edge et al. (2024).
Additionally, we explore the application of long LLMs in HawkBench, including vanilla LLMs, the prompt compression method Lingua-2 Pan et al. (2024), and long-context acceleration methods such as MInference Jiang et al. (2024). All baselines in the main experiments use Qwen2.5-7B-instruct as the generator Qwen et al. (2025), with BGE-M3 as the retriever Chen et al. (2023) and the top-$k$ set to 5 for all RAG methods.
For Level-1 and Level-2 tasks, which focus on factoid queries, we use Rouge-L and lexical F1-score as evaluation metrics. For Level-3 and Level-4 tasks, which involve rationale queries, we introduce a new evaluation metric, ${\mathcal{S}}$-F1, defined as:
${\mathcal{S}}\text{-F1}(A,A^{*})$ $=\frac{1}{2n}\sum_{i=1}^{n}\mathds{1}_{\{\text{LLM}(s_{i},A^{*})=\text{True}\}}$ (2)
$+\frac{1}{2n}\sum_{i=1}^{n}\mathds{1}_{\{\text{LLM}(s_{i}^{*},A)=\text{True}\}},$ (3)
where $A^{*}$ represents the ground-truth answer, and $A$ denotes the predicted answer. This metric evaluates: 1) The proportion of sentences in the ground-truth answer $s^{*}_{i}\in A^{*}$ that can be supported by the predicted answer. 2) The proportion of sentences in the predicted answer $s_{i}\in A$ that correctly reflect the meaning of the ground-truth answer.
The indicator function $\mathds{1}_{\text{condition}}$ returns 1 if the condition holds, and 0 otherwise.
Compared to lexical F1-score, ${\mathcal{S}}$-F1 evaluates sentence-level semantic equivalence between the ground-truth and predicted answers, making it a more robust metric for rationale-based tasks as lexical overlapping cannot infer the rationale equality. In addition to ${\mathcal{S}}$-F1, we also employ Rouge-L to evaluate Level-3 and Level-4 tasks.
3.2 Main Results
Resilience across Levels
Table 2 presents the performance of all baselines across the four task levels, averaged by domain. From these results, we draw several key insights:
(1) Standard RAG and Enhanced RAG methods perform well on factoid queries (Level-1 and Level-2), suggesting that these queries often rely on specific text spans that can be easily located with minimal reasoning or simple enhancements.
(2) Global RAG methods underperform on Level-1 and Level-2 tasks but excel on Level-3 and Level-4 tasks. This indicates that global reasoning is not beneficial for factoid queries and may even hinder performance. However, for rationale queries, which require synthesizing information from a broad range of text, global awareness helps gather more comprehensive evidence, leading to improved performance.
(3) Directly applying long LLMs to process the entire knowledge base is feasible but underperforms on factoid queries due to over-referencing and redundant noise. However, for rationale queries, long LLMs outperform vanilla RAG methods due to their strong reasoning ability over long contexts. Efficient long-context methods, such as accelerated pre-filling or prompt compression, yield performance comparable to vanilla LLMs.
@@ 3.2 Main Results
Figure 2 presents the experimental results across different levels and domains for selected methods. The results highlight how different methods perform across domain-specific knowledge:
(1) For structured knowledge sources, such as financial reports and legal documents, most methods perform well on factoid queries. The inherent clarity and precision of these texts reduce semantic ambiguity, improving retrieval accuracy.
(2) For explanatory texts, such as academic papers that focus on providing rationales, global RAG methods excel. Their global awareness enables them to effectively organize and integrate explicit reasoning from the knowledge base.
(3) For unstructured knowledge in domains like literature, art, and humanities—where texts contain higher semantic ambiguity—global RAG methods perform better on Level-4 tasks. This suggests that aggregating high-level implicit information is more effective for narrative-based content than for structured knowledge domains.
Figure 3: Evaluation performance across four levels for vanilla RAG with varying Top-$k$ selections.
@@
Figure 3 analyzes the impact of Top-k selection. The results show that while increasing Top-k introduces more knowledge into the generation process, it also increases redundancy. The trade-off between knowledge recall and precision varies across query levels. Factoid queries rely on precise evidence, and excessive redundancy significantly degrades performance. In contrast, rationale queries benefit from higher recall, as effective information aggregation requires a more comprehensive set of evidence from the knowledge base.
@@
1 0.6 1.0 29.1 20.9 1.7 (+$\infty$+ ∞)
2 0.7 2.0 32.7 21.5 2.0 (+$\infty$+ ∞)
3 1.6 2.1 48.3 33.4 3.0(+$\infty$+ ∞)
4 1.7 2.2 52.1 35.9 3.5(+$\infty$+ ∞)
Table 3: Task latency (queries per second) comparison across methods and levels. Experiments were conducted on an Nvidia A800-80G GPU using the Art dataset. GraphRAG employs GPT-4o for graph construction, which can take up to half an hour, denoted by +$\infty$+ ∞.
Table 3 presents a comparison of task latency across methods and task levels. The following insights can be drawn from the results: (1) Standard RAG methods are highly efficient, as the retrieval process is not sensitive to the size of the knowledge base. In contrast, long LLMs and global RAG methods experience a significant increase in latency across all tasks, while only improving performance on rationale tasks. (2) Long LLMs incur the highest latency for all task types but fail to deliver a clear performance advantage. This suggests that directly using the full knowledge base may not be a proper approach. (3) The graph construction process for GraphRAG relies heavily on robust model APIs, leading to substantial construction latency. However, once the graph is constructed, performance becomes efficient. This indicates that optimizing the process of perceiving the global knowledge base—such as accelerating the graph construction in GraphRAG or memory formation in MemoRAG—could be beneficial for improving performance on rationale queries.

### v2

@@ Abstract
In real-world information-seeking scenarios, users have dynamic and diverse needs, requiring RAG systems to demonstrate adaptable resilience. To comprehensively evaluate the resilience of current RAG methods, we introduce HawkBench, a human-labeled, multi-domain benchmark designed to rigorously assess RAG performance across categorized task types. By stratifying tasks based on information-seeking behaviors, HawkBench provides a systematic evaluation of how well RAG systems adapt to diverse user needs. Unlike existing benchmarks, which focus primarily on specific task types (mostly factoid queries) and rely on varying knowledge bases, HawkBench offers: (1) systematic task stratification to cover a broad range of query types, including both factoid and rationale queries, (2) integration of multi-domain corpora across all task types to mitigate corpus bias, and (3) rigorous annotation for high-quality evaluation. HawkBench includes 1,600 high-quality test samples, evenly distributed across domains and task types. Using this benchmark, we evaluate representative RAG methods, analyzing their performance in terms of answer quality and response latency. Our findings highlight the need for dynamic task strategies that integrate decision-making, query interpretation, and global knowledge understanding to improve RAG generalizability. We believe HawkBench serves as a pivotal benchmark for advancing the resilience of RAG methods and their ability to achieve general-purpose information seeking. We host our codes and data in this repository.
@@ 1 Introduction
Large Language Models (LLMs) excel in general reasoning and knowledge-based tasks but often struggle with timeliness and knowledge coverage gaps, particularly in specialized domains and user-specific data (OpenAI, 2023; DeepSeek-AI, 2024). To address these limitations, incorporating external knowledge has become a common approach, with Retrieval-Augmented Generation (RAG) emerging as an effective solution to enhance factual accuracy and adaptability (Zhu et al., 2024).
@@
During the information-seeking process using RAG, users may have a wide range of information needs, from simple factoid retrieval to more complex rationale-based queries (Qian et al., 2025; Zhao et al., 2024a). This versatility requires RAG systems to possess diverse capabilities, including accurate referencing and advanced reasoning skills.
@@
Recent advancements in RAG methods have enhanced vanilla RAG systems by targeting specific advanced capabilities. For instance, some methods focus on improving multi-hop reasoning to handle tasks with implicit information intents (Zhao et al., 2024b; Xu et al., 2024), while others address information aggregation tasks by constructing intermediate structures, such as graphs or memory modules, to better integrate relevant information (Qian et al., 2025; Edge et al., 2024).
@@
While these advancements enable RAG systems to effectively leverage external knowledge for specific tasks, their ability to generalize across diverse scenarios remains uncertain. A recent survey categorizes external knowledge-based tasks into distinct levels, emphasizing that no single method can effectively address all query types (Zhao et al., 2024a). This suggests that current RAG methods lack the resilience required for general-purpose information-seeking tasks, highlighting the need for a systematic evaluation of RAG methods across a broad range of information-seeking tasks, examining the resilience of these methods when faced with information-seeking tasks in any form.
@@
Existing public benchmarks for RAG evaluation focus narrowly on isolated dimensions of information-seeking tasks. For instance, LegalBench-RAG evaluates information-seeking tasks in the legal domain (Guha et al., 2023), MutiHop-RAG tests multi-hop reasoning (Tang and Yang, 2024), and CRAG emphasizes comprehensive evaluation on factual QA tasks (Yang et al., 2024). While these benchmarks excel in their targeted domains, they collectively fail to assess the resilience of RAG methods across stratified task types due to three critical limitations:
@@
First, fragmented evaluation protocols. Current benchmarks are siloed by design, each prioritizing distinct query types. This specialization creates inconsistent evaluation criteria, hindering fair comparisons of RAG performance across diverse task categories. Second, domain bias and knowledge leakage. Many benchmarks rely on heterogeneous knowledge bases (e.g., Wikipedia and web snippets), leading to corpus-dependent performance gaps that obscure true method capabilities. Worse, LLMs are often pretrained on these same sources (e.g., Wikipedia), inflating benchmark scores through memorization rather than genuine retrieval-augmented reasoning. Third, limited query diversity. Most benchmarks disproportionately emphasize factoid questions (e.g., "When was Einstein born?"), neglecting rationale-based queries (e.g., "Explain how relativity revolutionized physics") that require synthesis and contextual analysis. This narrow focus misaligns with real-world user needs, where information-seeking behaviors span both factual lookup and complex reasoning.
HawkBench is characterized by the following key features:
@@
Systematic Task Stratification – We systematically define four query types: (1) explicit factoid queries, (2) implicit factoid queries, (3) explicit rationale queries, and (4) implicit rationale queries. This stratification, inspired by Zhao et al. (2024a) with refined modifications, ensures comprehensive task coverage. Importantly, all query types share the same underlying knowledge distribution, allowing for direct and fair performance comparisons across different tasks.
@@ 2.3 Comparison of the Four Query Levels
In Figure 1 (right), we compare the four query levels across two aspects: Reference and Reasoning. First, in terms of Reference, the amount of relevant knowledge required increases from Level 1 to Level 4 queries, reflected in the mutual information between the knowledge base and retrieved knowledge, $I({\mathcal{X}},{\mathcal{Z}})$. Level 1 queries require minimal knowledge, as answers are directly extractable from a few text chunks. In contrast, higher-level queries, such as Level 3 and Level 4, require synthesizing information from a broader range of texts. Second, in terms of Reasoning, complexity increases across levels due to the growing semantic gap between retrieved knowledge and the final answer. For Level 1 queries, reasoning is minimal, but for Level 3 and Level 4 queries, more reasoning is needed to connect multiple, loosely connected pieces of information. This is reflected in the decreasing mutual information $I({\mathcal{Z}},{\mathcal{Y}})$ as redundant information is filtered out during refinement.
@@ 2.4 Construction
While most current LLMs are proficient in general world knowledge due to their training on large-scale corpora, they often lack coverage in specialized, domain-specific areas. To address this gap, HawkBench incorporates 229 domain-specific texts into its knowledge base. These texts are carefully selected from a larger collection of long texts gathered across diverse domains, which can also serve as a global corpus for retrieval. The selected 229 contexts span a wide range of domains, including professional textbooks (manually labeled into categories such as technology, humanities, art, and science), financial reports, legal contracts, novels, and academic papers. This diverse and comprehensive collection ensures that HawkBench can thoroughly evaluate the domain resilience of RAG methods by covering a broad range of user information needs.
@@
(2) Question-Answer Pair Generation: The system prompts the LLM agent using built-in QA generation prompts to produce initial question-answer pairs. During this step, the system first samples from the knowledge base, selecting a random text span of varying lengths based on the task type. For Level-1 tasks, approximately 1K tokens are used as the context. For Level-2 tasks, we use a retrieval system retrieves the top-10 passages based on the generated L1 query, selecting five passages to prompt the agent to transform explicit factoid queries into implicit intent queries. For Level-3 and Level-4 tasks, up to 120K tokens are sampled as the knowledge context to guide the agent in generating information aggregation queries, with different prompts controlling the process. The codes for annotation system and all built-in prompts are in this repository.
L Discard % Edit % Ave. Time Total Time
@@
Figure 2: Statistical Details of Construction.
@@
3 Experiment
3.1 Baselines and Metrics
To investigate the resilience of RAG methods on HawkBench, we select the following representative baseline methods: Vanilla RAG: This method retrieves the top passages as context. Enhanced RAG Methods: HyDE (Gao et al., 2023) generates a hypothetical document to enhance query retrieval. RQRAG (Chan et al., 2024) rewrites the input query into sub-queries to refine retrieval. Global RAG: These methods index the knowledge base into an intermediate form to enhance global awareness. This includes memory-based methods such as MemoRAG (Qian et al., 2025) and graph-based methods like GraphRAG (Edge et al., 2024).
Additionally, we explore the application of long LLMs in HawkBench, including vanilla LLMs, the prompt compression method Lingua-2 (Pan et al., 2024), and long-context acceleration methods such as MInference (Jiang et al., 2024). All baselines in the main experiments use Qwen2.5-7B-instruct as the generator (Team, 2025), with BGE-M3 as the retriever (Chen et al., 2023) and the top-$k$ set to 5 for all RAG methods.
For Level 1 and Level 2 tasks, which focus on factoid queries, we use Rouge-L and lexical F1-score as evaluation metrics. These metrics emphasize surface-form lexical overlap and are well suited for evaluating fact-based answers.
For Level 3 and Level 4 tasks, which involve rationale queries, we introduce a new evaluation metric, denoted as $\mathcal{S}$-F1, to robustly assess sentence-level semantic equivalence between the ground-truth and predicted answers. Specifically, let $A^{*}$ denote the ground-truth answer and $A$ the predicted answer. We tokenize both $A$ and $A^{*}$ into sentences $\{s_{i}\}$ and $\{s^{*}_{i}\}$, respectively. Then, $\mathcal{S}$-F1 is defined as:
$\mathcal{S}\text{-F1}(A,A^{*})=\frac{1}{2n}\sum_{i=1}^{n}\mathds{1}_{\{\text{LLM}(s_{i},A^{*})=\text{True}\}}+\frac{1}{2m}\sum_{i=1}^{m}\mathds{1}_{\{\text{LLM}(s^{*}_{i},A)=\text{True}\}},$ (2)
where $\mathds{1}_{\text{condition}}$ is an indicator function that returns $1$ if the condition holds and 0 otherwise, $n$ is the number of sentences in $A$, and $m$ is the number of sentences in $A^{*}$.
Intuitively, $\mathcal{S}$-F1 computes the average of: (1) Precision: the proportion of sentences in the predicted answer $s_{i}\in A$ that are judged by a strong LLM to be semantically supported by the ground-truth answer $A^{*}$. (2) Recall: the proportion of sentences in the ground-truth answer $s^{*}_{i}\in A^{*}$ that are judged to be semantically supported by the predicted answer $A$.
Here, “supported by” means that for a given sentence, the LLM judge determines whether its meaning or rationale is present, possibly rephrased but semantically equivalent, in the other answer. More concretely, we apply the following process: (1) Each predicted answer is split into sentences. For each $s_{i}\in A$, the LLM judge is prompted to return a binary decision (0/1) indicating whether the content of $s_{i}$ is covered by $A^{*}$. (2) Each ground-truth answer is similarly split into sentences, and for each $s^{*}_{i}\in A^{*}$, we query whether its rationale is reflected in $A$.
Compared to lexical F1-score, $\mathcal{S}$-F1 moves beyond surface-form matching and directly evaluates sentence-level semantic alignment between $A$ and $A^{*}$, making it a more robust metric for rationale-based tasks where lexical overlap alone cannot capture equivalence. For completeness, we also report Rouge-L scores alongside $\mathcal{S}$-F1 when evaluating Level 3 and Level 4 tasks.
3.2 Main Results
Resilience across Levels
Table 1 presents the performance of all baselines across the four task levels, averaged by domain. From these results, we draw several key insights: (1) Standard RAG and Enhanced RAG methods perform well on factoid queries (Level-1 and Level-2), suggesting that these queries often rely on specific text spans that can be easily located with minimal reasoning or simple enhancements. (2) Global RAG methods underperform on Level-1 and Level-2 tasks but excel on Level-3 and Level-4 tasks. This indicates that global reasoning is not beneficial for factoid queries and may even hinder performance. However, for rationale queries, which require synthesizing information from a broad range of text, global awareness helps gather more comprehensive evidence, leading to improved performance. (3) Directly applying long LLMs to process the entire knowledge base is feasible but underperforms on factoid queries due to over-referencing and redundant noise. However, for rationale queries, long LLMs outperform vanilla RAG methods due to their strong reasoning ability over long contexts. Efficient long-context methods, such as accelerated pre-filling or prompt compression, yield performance comparable to vanilla LLMs.
@@ 3.2 Main Results
Figure 3 presents the experimental results across different levels and domains for selected methods. The results highlight how different methods perform across domain-specific knowledge: (1) For structured knowledge sources, such as financial reports and legal documents, most methods perform well on factoid queries. The inherent clarity and precision of these texts reduce semantic ambiguity, improving retrieval accuracy.
Figure 4: Evaluation performance across four levels for RAG with varying Top-$k$ selections.
(2) For explanatory texts, such as academic papers that focus on providing rationales, global RAG methods excel. Their global awareness enables them to effectively organize and integrate explicit reasoning from the knowledge base. (3) For unstructured knowledge in domains like literature, art, and humanities—where texts contain higher semantic ambiguity—global RAG methods perform better on Level-4 tasks. This suggests that aggregating high-level implicit information is more effective for narrative-based content than for structured knowledge domains.
@@
Figure 4 systematically investigates the impact of Top-k selection using vanilla RAG. The results show that while increasing Top-k introduces more knowledge into the generation process, it also increases redundancy. The trade-off between knowledge recall and precision varies across query levels. Factoid queries rely on precise evidence, and excessive redundancy significantly degrades performance. In contrast, rationale queries benefit from higher recall, as effective information aggregation requires a more comprehensive set of evidence from the knowledge base.
@@
Table 2 presents a comparison of task latency across methods and task levels. The following insights can be drawn from the results:
(1) Standard RAG methods are highly efficient, as the retrieval process is not sensitive to the size of the knowledge base. In contrast, long LLMs and global RAG methods experience a notable increase
@@
1 0.6 1.0 29.1 20.9 1.7 ($\infty$)
2 0.7 2.0 32.7 21.5 2.0 ($\infty$)
3 1.6 2.1 48.3 33.4 3.0($\infty$)
4 1.7 2.2 52.1 35.9 3.5($\infty$)
Table 2: Task latency (queries per second) comparison across methods and levels. Experiments were conducted on an Nvidia A800-80G GPU using the Art dataset. GraphRAG employs GPT-4o for graph construction, which can take up to half an hour, denoted by $\infty$.
in latency across all tasks, while only improving performance on rationale tasks. (2) Long LLMs incur the highest latency for all task types but fail to deliver a clear performance advantage. This suggests that directly using the full knowledge base may not be a proper approach. (3) The graph construction process for GraphRAG relies heavily on robust model APIs, leading to substantial construction latency. However, once the graph is constructed, performance becomes efficient. This indicates that optimizing the process of perceiving the global knowledge base—such as accelerating the graph construction in GraphRAG or memory formation in MemoRAG—could be beneficial for improving performance on rationale queries.
Retrieval Strategy Analysis
In addition to comparing different RAG architectures, we further investigate the impact of retrieval strategies on performance across various task levels. Specifically, we evaluate three types of retrievers: dense retrieval, sparse retrieval, and a hybrid approach that combines both. The goal is to understand how the choice of retriever influences the resilience and adaptability of RAG systems under different information-seeking challenges.
Figure 5: Performance of different retrieval strategies across the four HawkBench levels.
Figure 5 presents the performance of representative RAG methods (vanilla RAG, RQRAG, and MemoRAG) using each retrieval strategy across the four task levels in HawkBench. The results demonstrate that retrieval strategy has a substantial impact on downstream performance. Dense and hybrid retrievers consistently outperform sparse retrievers, particularly on rationale-intensive tasks (Levels 3 and 4), where retrieving semantically rich information is crucial. Notably, methods that incorporate additional retrieval cues—such as query rewriting in RQRAG or memory-guided retrieval in MemoRAG—benefit significantly from hybrid retrieval. This suggests that hybrid retrievers enhance the likelihood of capturing diverse, relevant evidence when guided by auxiliary signals. These findings underscore the importance of retrieval design in RAG pipelines, especially when targeting general-purpose or reasoning-intensive tasks. Future research may explore adaptive retrieval modules that dynamically select the most suitable retrieval strategy based on task characteristics.